# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### v2

@@ 3.3 Dataset Statistics
Furthermore, Figure 6 and Table 3 (Supplementary Material) display the frequency distribution of the instrument instances per class, also demonstrating a natural long-tail effect. The reason for this distribution can also be observed in the co-occurrence matrices detailed in the Supplementary Material. The Monopolar Curved Scissors (MCS) and Bipolar Forceps (BF) are the most frequent instruments. This over-representation is primarily because these tools are the most actively used in robot-assisted surgeries, owing to their versatility and capacity to perform most surgical actions. They are also among the most crucial robotic instruments of the da Vinci Surgical System (dVSS). Following these, the Large Needle Driver (LND), another typical robotic instrument of the dVSS, also presents high frequencies as it is utilized for all suturing stages, but it has reduced functionalities compared to the MCS and BF. While also commonly used, the Prograsp Forceps (PF) and the Suction Instrument (SI) have more specific roles in surgery and often function passively and steadily. The latter is additional to the dVSS. Finally, instruments like the Clip Applier (CA) and other Laparoscopic Instruments (LI) are also additional to the dVSS and employed for specific purposes during certain stages of the surgery.
Figure 7 and Table 4 (Supplementary Material) illustrate the distribution of the number of instrument instances per video frame. We observe that each video frame contains 1 to 5 instrument instances (2.23 average of instrument instances per frame), with approximately 80% of frames containing between 2 and 3 instances. These statistics are characteristic of robotic surgeries, where surgeons mainly use two robotic instruments simultaneously with their hands. The additional presence of non-robotic instruments (SI, LI, and CA) manipulated by an assistant or a third robotic instrument left stationary increases the possible number of instances per frame and contributes to the diversity in surgical visual scenes. Situations involving four or five instruments are less common and usually indicate specific complex scenarios within the surgery. The upper limit of five instruments per frame is due to the inherent limitations of the dVSS and the specific complexities associated with Radical Prostatectomies.
Fig. 7: Distribution of the percentage of frames presenting (n) instrument instances per frame on each data split. The percentage of frames at 35 fps that have (n) instrument instances in it. The figure is best viewed in color.
Finally, Figure 9 and Table 5 (Supplementary Material) present the frequency and normalized distribution of the atomic action instances per class. Again, the distribution of atomic actions suggests a long-tail behavior where still (11), hold (5), and travel (13) are the most frequently occurring actions in the dataset. Indeed, during the surgery, most instruments are either used to hold and maintain tissues and objects in place or to move these tissues and objects to a different location. Later on, Push (9) and Pull (8) represent more specialized movements and thus are less present. Suction (12) is an action that can only be performed by the SI (as shown in the instrument-action co-occurrence matrix in Figure 14 in the Supplementary Material). Nonetheless, when on the frame, the SI very frequently performs this action, which explains the higher frequency of this action. The remaining actions present lower frequencies as they are much more specialized and usually done by a specific type of instrument in shorter time frames. For instance, Cauterize (1) can only be performed by the BF and MCS, and Cut (3) is only performed by the MCS. Similarly, Open (6) can be done by all the instruments with tips (all except SI) but for very short durations. Lastly, the Other class (14) is the most scarce category since it is a composition of Staple and Wash, which can only be performed by the CA and the SI, respectively, and only happen in very particular moments of the surgery.
@@
The statistical distributions presented in Figures 23678 and 9 demonstrate that our dataset maintains similar distributions across all the predefined dataset splits. First, this consistency highlights the reliability of our entire dataset and its predefined cross-validation folds and test set, showing that all the categories in all the proposed tasks keep comparable frequency distributions in all splits. Second, this observation validates our annotation process as it proves the efficiency of our designed annotation protocols and our multiple stages of data validation and correction.
We include further statistical analysis in C of the Supplementary Material. This section provides information about the total number of instances in each task of the GraSP dataset, the transition probabilities for both phase and steps, the intra-task categories co-occurrences for the instruments and actions tasks, and the detailed inter-task categories co-occurrences for instruments-phases, instrument-steps, and instruments-actions in GraSP. Our previously presented statistics and the additional statistics in the Supplementary Material prove the intrinsic complementarity among the proposed tasks in GraSP. We previously established the relation between phases as steps where their per-class frequency and duration distributions demonstrated similar patterns related to the semantic meaning of the categories. Regarding this, the most durable phase categories encompass many of the most durable step categories, and the least durable phases encompass many short steps. Likewise, there is a clear relation between the surgical instrument and action categories, where many actions can be performed only by specific instruments. Thus, the frequency of different instruments affects the frequencies of actions. Additionally, the co-occurrence matrixes demonstrate an evident correspondence between instruments and phases or steps as, intuitively, different instruments are more present or completely absent in different phases and steps.
@@ 3.4 Evaluation Metrics
To evaluate the performance of phase and step recognition, we keep the originally proposed metric in Valderrama et al. (2022): the mean Average Precision ($mAP$) per frame. We chose this metric as it is a standard and robust metric commonly used for frame-wise evaluation of action recognition (Caba Heilbron et al. (2015)). We use the implementation of Python’s Scikit-learn library 555sklearn.metrics.average_precision_score for the calculation of this metric.
Regarding the instrument segmentation task, we adopt the conventional instance-based mean Average Precision ($mAP$) from PASCAL VOC (Everingham et al. (2015)) to exploit the instance-wise nature of this task entirely. Following Everingham et al. (2015), we adopt the $mAP@0.5IoU_{segm}$ metric, which uses a minimum Intersection over Union (IoU) threshold of 50% between the ground-truth masks and the predicted masks to categorize predictions into true positive predictions. We hope that including the $mAP@0.5IoU_{segm}$ will encourage the community to approach the surgical instrument segmentation problem from a similar evaluation perspective. We calculate the $mAP@0.5IoU_{segm}$ metric using a modification of Activity Net’s (Caba Heilbron et al. (2015)) implementation  [^6] https://github.com/activitynet/ActivityNet.
Additionally, we maintain the usual semantic segmentation evaluation metrics we introduced in González et al. (2020). These metrics include the Mean Intersection over Union (mIoU), Intersection over Union (IoU), and Mean Class Intersection over Union (mcIoU). We kept these metrics as they are highly stringent metrics that have been the standard evaluation methodology for the instrument segmentation task in multiple works. We utilize the implementation provided in Ayobi et al. (2023)  [^7] https://github.com/BCV-Uniandes/MATIS.
Lastly, for the Atomic Action Detection task, we adhere to the original evaluation framework from AVA (Gu et al. (2018); Li et al. (2020)) and Activity Net (Caba Heilbron et al. (2015)), which is, in turn, based on PASCAL VOC’s object detection evaluation (Everingham et al. (2015)). As previously stated, the atomic actions in our framework are performed by an instrument and, as such, are centered on it. Thus, locating surgical atomic visual actions consists of locating the acting instrument that performs the actions. However, the primary objective is to accurately detect the presence and location of these actions rather than delineating the exact shape of the agent that executes them. Hence, an object detection approach is better than an object segmentation approach for the atomic action location task.
For this reason, regardless of our instrument segmentation annotations, we maintain the original $mAP@0.5IoU_{box}$ metric from AVA, which is analogous to the previously explained $mAP@0.5IoU_{segm}$ metric but using the ground truth and predicted bounding boxes to calculate the IoU. Finally, due to the multi-target fashion of atomic action detection, where each instrument instance can perform up to three actions, we independently evaluate the presence of each atomic action in an instrument instance instead of assessing whether the entire expected combination of ground truth actions was predicted. This methodology is a more robust and sensitive way of measuring the atomic action detection performance. Once again, we use Activity Net’s implementation of the $mAP@0.5IoU_{box}$ metric for atomic visual actions detection.
@@ 4 Transformers for Actions, Phases, Steps and Instrument Segmentation
Fig. 10: TAPIS utilizes Mask2Former (Cheng et al. (2022)) as its instrument segmentation baseline and MViT (Fan et al. (2021)) as its video feature extractor. On the one hand, the segmentation baseline takes an input frame and computes a set of per-segment embeddings by cross-attending learnable object queries with multi-scale features extracted with a visual backbone and processed with a pixel decoder. Then, the model generates a set of binary masks with the dot multiplication between the linearly transformed per-segment embedding with pixel features from a pixel decoder. On the other, the video feature extractor uses a time window centered on the keyframe used for segmentation and computes a sequence of spatio-temporal embeddings by patchifying the input video and applying multiple stages made by blocks of Multi Head Pooling Attention. These stages progressively reduce the sequence length and increase the number of features. For the long-term tasks, the frame classification head linearly classifies the learnable class token of the video feature extractor into either a phase or a step category. For the short-term tasks, a region classification head cross-attends the precalculated segment embeddings from the instrument segmentation with the spatio-temporal features, and then they are linearly classified into instruments or action classes. The figure is best viewed in color.
@@ 4.1 Instrument Segmentation Baseline
We follow MATIS (Ayobi et al. (2023)), our previous approach for the instrument segmentation task, which adapts our winning solution of the SAR-RARP50 challenge 888synapse.sar-rarp50_challenge. Hence, we employ Mask2Former (Cheng et al. (2022)) as our primary instrument segmentation baseline and region proposal method. We select this method as it is highly efficient and has demonstrated excellent performance in instance segmentation of natural and robot-assisted surgery images. In this section, we briefly review Mask2Former’s methodologies.
The general intuition behind Mask2Former is to generate a fixed set of $\mathcal{N}$ binary instance masks with their corresponding class probability distribution. For this purpose, Mask2Former adapts DETR’s (Carion et al. (2020)) and MaskFormer’s (Cheng et al. (2021)) meta architectures based on processing a set of $\mathcal{N}$ learnable object queries into $\mathcal{N}$ per-segment embedding by using a transformer decoder that cross-attends the learnable object queries with the image features extracted from the backbone. However, Mask2Former utilizes a Multi-Scale Deformable Attention (Zhu et al. (2021)) pixel decoder with upsampling layers to compute multi-scale per pixel features to feed as image features for the transformer decoder. Also, it replaces the conventional global cross-attention with masked attention, which employs binary masks generated from the pixel embeddings and the transformed object queries to restrict attention and focus on relevant regions. These enhancements significantly increase training efficiency and boost segmentation performance. Finally, a linear classifier projects each outputted segment embedding into a class probability, and simultaneously, a multi-layer perceptron (MLP) transforms the segment embeddings into mask embeddings to generate the final binary masks by the dot multiplication with the per-pixel features. Figure 10 displays this methodology at the top.
We refer the reader to the original Mask2Former (Cheng et al. (2022)) and MATIS (Ayobi et al. (2023)) papers for further details on the Mask2Former architecture and its adaptation for surgical instrument segmentation. Furthermore, our TAPIS model utilizes Mask2Former’s region proposals two-fold. First, as in MATIS, to improve temporal consistency in instrument segmentation by reclassifying region proposals utilizing temporal context provided by a video feature extractor. Second, to handle the implicit need for instrument localization in atomic action detection. Thereby, TAPIS uses the corresponding segment embeddings of the proposed regions as inputs for a region classification head that enhances the segment classification process or predicts the atomic actions performed by the proposed instruments. We provide further insights about our region classification module in Section 4.3.
@@ 4.2 Video Feature Extractor
Following our previous work, we employ MViT (Fan et al. (2021)) as our video feature extractor to capture intricate details across various space-time scales by leveraging transformers. Once again, we briefly review MViT and refer the reader to the original paper for further details (Fan et al. (2021)).
MViT, an extension of ViT Dosovitskiy et al. (2021) for video-level recognition, is based on consecutive stages of multiple transformer blocks. This architecture first partitions the image into overlapping patches, and then, through each stage, it contracts the spatio-temporal dimensions (sequence length/resolution) while the channel dimensions (features length) expand. To achieve this, MViT replaces the Multi-Head Attention employed in ViT with Multi-Head Pooling Attention (MHPA). This adaptation involves pooling the queries, keys, and values at the beginning of each stage to reduce the spatio-temporal resolution before computing attention. Additionally, MViT uses linear projections at the end of each stage to increase the channel dimension for the following stages. This process creates a feature pyramid network, with the initial stages operating at more detailed resolutions and minor features and the final stages processing shorter sequences with more complex spatio-temporal features. Additionally, MViT follows BERT’s (Devlin et al. (2019)) and ViT’s (Dosovitskiy et al. (2021)) methodology of including a learnable class token that acts as a single embedding of the entire sequence for classification tasks.
We utilize the classification nature of MViT for phases and steps recognition by classifying an entire time window centered on a keyframe with the class of the middle frame. For this purpose, we use a frame classification head that employs a linear classifier on the output class token of the video features extractor to derive the final Phase or Step class probability distribution. Figure 4.4 portrays the video feature extractor and the frame classification head at the bottom. The following section extends how we use the remaining embeddings for region classification and action recognition.
@@ 4.3 Region Classification Head
We introduce a novel region classification head for instrument classification and atomic action detection. Our region classification head builds upon the box classification head proposed in Valderrama et al. (2022), which mainly combines precalculated region features from a frame-wise box proposal network with the sequence of temporal features generated by the video feature extractor. In this work, we leverage transformers’ cross-attention to correlate region and temporal features. For this purpose, we input the per-segment embeddings of the proposed regions into a cross-attention layer that performs Multi-Head Attention over the entire sequence of spatio-temporal features from the video feature extractor. Consequently, this layer can fully exploit the spatio-temporal resolution to enrich the region features with relevant temporal context. Finally, we use a linear classification layer to project the enriched region features outputted by the cross-attention layer into either instrument or atomic action class probability distributions.
Although the region classification head can be agnostic to the specific region proposal method, we use our instrument segmentation baseline (explained in Section 4.1) to obtain pixel-wise instrument instance region proposals and their per-segment embeddings. Hence, our region classification head cross-attends these shape-wise embeddings with the temporal cues of the video feature extractor to enhance the instrument region classification process and predict each instrument’s atomic actions. Figure 10 shows this head in the upper-right part.
Table 4: Performance of TAPIS in GraSP. We report the results of TAPIS on the test set of GraSP, and we compare its performance with a CNN-based baseline (SlowFast), the TAPIR model proposed in Valderrama et al. (2022) and a Video Swin Transformer-based implementation (TAPIS-VST). We present the number of parameters (Params.) and floating point operations (FLOPs) of each model without counting the region proposal network. The best results are shown in bold.
Method FLOPs (G) Params. (M) Long-Term Tasks $mAP$ (%) Short-Term Tasks $mAP@0.5IoU_{box}$ (%) Instrument Segmentation Metrics (%)
Phases Steps Instruments Atomic Actions $mAP@0.5IoU_{segm}$ mIoU IoU mcIoU
SlowFast 81 33 70.70 46.23 74.33 22.01 71.32 77.16 72.26 58.75
TAPIR 71 36 74.55 49.42 74.43 25.57
TAPIS-VST 66 96 70.88 46.99 90.29 33.10 89.58 86.36 83.51 77.54
TAPIS 71 44 76.07 49.42 89.85 39.26 89.10 86.61 83.38 77.42
@@ 4.4 Implementation Details
Instrument Segmentation Baseline: We used the official PyTorch implementation of Mask2Former [^9] https://github.com/facebookresearch/Mask2Former. We trained Mask2Former for 100 epochs on 4 NVIDIA Quadro RTX 8000 GPUs using a batch size of 32 for a ResNet50 backbone (He et al. (2016)) and 4 for a Swin Large (Liu et al. (2021)) backbone. We used a base learning rate of $1e^{-5}$ decayed by 0.1 in epochs 50 and 75 and an ADAMW optimizer (Loshchilov and Hutter (2019)). We used the publicly available pretrained weights for instance segmentation on the MS-COCO dataset (Lin et al. (2014)). We selected the top 5 scoring regions with more than 0.1 confidence scores for inference and region selection.
Phase and Step Recognition: We used MViT’s official implementation  [^10] https://github.com/facebookresearch/SlowFast. Using a cross-entropy loss, we trained MViT with its Base configuration for 30 epochs on a single NVIDIA Quadro RTX 8000 GPU. We trained our models using MViT’s pretrained weights in Kinetics 400 (Kay et al. (2017)). We used windows of 16 frames with a stride of 4 seconds (0.25fps sampling rate) for phases recognition and 1 second (1fps sampling rate) for steps recognition, and we used a batch size of 24 for both tasks. We used a cosine learning rate scheduler (Loshchilov and Hutter (2017)) with 5 warm-up epochs and an SGD optimizer with a base learning rate of $31.25e-$ 3 and an end learning rate of $41e-$ 4.
Instrument Region Classification: We trained TAPIS for 20 epochs in a single NVIDIA Quadro RTX 8000 GPU with a batch size of 24. We used time windows of 16 frames with a stride of 0.5 seconds (2fps sampling rate). We used a cosine learning rate (Loshchilov and Hutter (2017)) with 0 warm-up epochs and an SGD optimizer with a base learning rate of $21.25e-$ 2 and an end learning rate of $41e-$ 4. We trained our models using the pretrained weights in phases recognition.
Atomic Action Detection: We trained TAPIS for 50 epochs in a single NVIDIA Quadro RTX 8000 GPU with a batch size of 24 and binary a cross-entropy loss. We used the pretrained weights in phases recognition as initialization. We used time windows of 16 frames with the smallest stride possible (0.033 stride or 30fps sampling rate). We used a cosine learning rate (Loshchilov and Hutter (2017)) with 0 warm-up epochs and an SGD optimizer with a base learning rate of $11.25e-$ 1 and an end learning rate of $21e-$ 2.
Benchmarking: We performed all our optimization and design experiments in the defined cross-validation folds. For testing, we re-trained our best models selected during cross-validation in the entire training set (fold one plus fold two) and tested them on the defined test set. We do this to take full advantage of the entire extension of the training data. Since we maintain the atomic action evaluation as a detection evaluation, we consider the bounding boxes generated directly from the edges of the predicted instrument masks as bounding box predictions (no further bounding box coordinates regression). Finally, our collected videos include multiple flags, signals, and indicators about the mounted instruments and their usage. To avoid biases and correctly assess our models’ recognition capabilities, we removed all these signals from the sampled images throughout all our models and experiments. We provide an example of this frame modification procedure in the Supplementary Material.

### v3

@@ 3.4 Dataset Statistics
Fig. 8: Distribution of the percentage of frames presenting (n) instrument instances per frame on each data split. The percentage of frames at 35-second intervals with (n) instrument instances. The figure is best viewed in color.
Figure 9 and Tables 5 and 6 (Supplementary Material) present the frequency distribution of the atomic action instances per class. Again, the distribution of atomic actions suggests a long-tail behavior where still, hold, and travel are the most frequent since most instruments are used to hold and maintain tissues and objects in place or to move them to a different location. On the contrary, Push, Pull, and Suction represent more specialized movements, and the remaining actions like Cauterize, Cut, Open and Other, are usually done by a specific type of instrument in very short time windows, thus making them the least frequent.
@@
Finally, the statistical distributions presented in Figures 3478 and 9 demonstrate that our dataset maintains similar distributions across all the predefined dataset splits, highlighting the reliability of our entire dataset and annotation processes. Additionally, we include further statistical analysis in C of the Supplementary Material. This section provides information about the size (Figure 6) and total number of instrument instances (Table 4) in each fold of the GraSP dataset, and the co-occurrences matrixes for instruments-phases (Figure 14), instrument-steps (Figure 15), and instruments-actions (Figure 16).
Our dataset statistics prove the intrinsic complementarity among the proposed tasks in GraSP. For instance, the most time-consuming phase categories encompass many of the most time-consuming step categories and vice versa. Likewise, multiple actions can only be performed by specific instruments, thus causing a correlation between instrument and action frequencies. Additionally, the co-occurrence matrixes demonstrate an evident correspondence between instruments and phases or steps as different instruments are more present or completely absent in different surgical stages.
@@ 3.5 Evaluation Metrics
For phase and step recognition, we keep the mean Average Precision ($mAP$) originally proposed in [84] as it is an accepted standard for frame-wise evaluation of action recognition ([10]). Additionally, we include the video-wise $F1score$ as suggested in [27]555sklearn.metrics Following standard practice ([83, 27]) we calculate these metrics on frames sampled at 1fps.
For the instrument segmentation task, we adopt the conventional instance-based mean Average Precision ($mAP@0.5IoU_{segm}$) from PASCAL VOC ([23]) to promote research toward instance-based evaluation of this task. [^6] https://github.com/activitynet/ActivityNet Additionally, we maintain the standard and stringent semantic segmentation metrics we introduced in [32], which include the Mean Intersection over Union (mIoU), Intersection over Union (IoU), and Mean Class Intersection over Union (mcIoU). [^7] https://github.com/BCV-Uniandes/MATIS.
For the Atomic Action Detection task, we adhere to the evaluation framework established by AVA ([34, 50, 10]). Since surgical atomic actions are instrument-centric and the primary goal is to locate these actions rather than delineating the exact shape of the executing agent, we maintain the original object detection metric ($mAP@0.5IoU_{box}$) from AVA, applied to instrument bounding boxes instead of segmentation masks. Finally, given the multi-target nature of atomic action detection, where each instrument instance can perform up to three actions simultaneously, we evaluate the prediction of each atomic action category independently for each instrument instance instead of assessing the entire expected combination of ground truth actions. This methodology offers a more robust and sensitive way of measuring the atomic action detection performance.
@@ 4 Transformers for Actions, Phases, Steps and Instrument Segmentation
Fig. 10: TAPIS utilizes Mask2Former ([12]) as its instrument segmentation baseline and MViT ([24]) as its video feature extractor. On the one hand, the segmentation baseline takes an input frame and computes a set of per-segment embeddings by cross-attending learnable object queries with multi-scale features extracted with a visual backbone and processed with a pixel decoder. Then, the model generates a set of binary masks with the dot multiplication between the linearly transformed per-segment embedding with pixel features from a pixel decoder. On the other, the video feature extractor uses a time window centered on the keyframe used for segmentation and computes a sequence of spatio-temporal embeddings by patchifying the input video and applying multiple stages made by blocks of Multi Head Pooling Attention. These stages progressively reduce the sequence length and increase the number of features. For the long-term tasks, the frame classification head linearly classifies the learnable class token of the video feature extractor into either a phase or a step category. For the short-term tasks, a region classification head cross-attends the precalculated segment embeddings from the instrument segmentation with the spatio-temporal features, and then they are linearly classified into instruments or action classes. The figure is best viewed in color.
@@ 4.1 Instrument Segmentation Baseline
We follow MATIS’s aproach ([5]) and adopt Mask2Former ([12]) as our primary instrument segmentation baseline and region proposal method. Generally, Mask2Former employs a transformer decoder that cross-attends a set of $\mathcal{N}$ learnable object queries with image features extracted from a backbone, therefore transforming these queries into $\mathcal{N}$ per-segment embeddings. Additionally, Mask2Former uses a Multi-Scale Deformable Attention pixel decoder ([94]) with upsampling layers to compute multi-scale per-pixel features to feed into the transformer decoder. Mask2Former replaces conventional global cross-attention with masked attention, restricting attention to relevant regions using binary masks generated from the pixel embeddings and the transformed object queries. The output segment embeddings are then projected into class probabilities through a linear classifier, while a multi-layer perceptron (MLP) transforms the embeddings into mask embeddings. These mask embeddings are combined with the per-pixel features via dot multiplication to generate the final binary masks. Figure 10 illustrates this methodology at the top.
We refer the reader to the original Mask2Former ([12]) and MATIS ([5]) papers for further details on the Mask2Former architecture and its adaptation for surgical instrument segmentation. We explain how TAPIS utilizes Mask2Former’s regions for atomic action detection and instrument reclassification in Section 4.3. Finally, we follow [5] and refer to our instrument segmentation baselines as TAPIS Frame in our results tables.
@@ 4.2 Video Feature Extractor
Following our previous work, we employ MViT ([24]) as our video feature extractor to capture intricate details across various space-time scales by leveraging transformers. Once again, we briefly review MViT and refer the reader to the original paper for further details ([24]).
MViT is based on consecutive stages of transformer blocks. This architecture first partitions the image into overlapping patches and, through each stage, contracts the spatio-temporal dimensions (sequence length/resolution) while expanding the channel dimensions (features length). To achieve this, MViT uses Multi-Head Pooling Attention (MHPA) that pools queries, keys, and values at the beginning of each stage to reduce the spatio-temporal resolution before computing attention, while a linear projection at the end of each stage increases the channel dimension. This process creates an internal feature pyramid network, with the initial stages operating at more detailed resolutions with minor features, and the final stages processing shorter sequences with complex spatio-temporal features. Additionally, MViT includes a learnable class token to act as a single embedding of the entire sequence for classification tasks.
We utilize MViT’s classification nature for phase and step recognition by classifying an entire time window with the class of the middle frame. For this purpose, we use a frame classification head that employs a linear classifier on the output class token of the video features extractor to derive the final Phase or Step class probability distribution. Figure 4.4 portrays the video feature extractor and the frame classification head at the bottom. The following section extends how we use the remaining embeddings for region classification and action recognition.
Table 4: Performance of TAPIS in GraSP. We report the results of TAPIS on the test set of GraSP, and we compare its performance with a CNN-based baseline (SlowFast), the TAPIR model proposed in [84] and a Video Swin Transformer-based implementation (TAPIS-VST). We present the number of parameters (Params.) and floating point operations (FLOPs) of each model without counting the region proposal network. The best results are shown in bold.
Method FLOPS (G) Params (M) Phases Steps Atomic Actions Instrument Segmentation
$mAP$ F1 score $mAP$ F1 score $mAP@0.5IoU_{box}$ $mAP@0.5IoU_{box}$ $mAP@0.5IoU_{segm}$ mIoU IoU mcIoU
SlowFast 81 33 68.77 54.70 46.35 37.08 22.01 74.33 71.32 77.16 72.26 58.75
TAPIR 71 36 72.59 58.62 50.24 43.82 25.57 74.43
TAPIS-VST 66 96 76.24 60.80 50.74 41.94 33.10 90.29 89.58 86.36 83.51 77.54
TAPIS 71 44 76.72 63.42 52.01 45.78 39.26 89.85 89.10 86.61 83.38 77.42
@@ 4.3 Region Classification Head
We introduce a novel region classification head for instrument classification and atomic action detection. Our region classification head builds upon the box classification head proposed in [84]. In this work, we leverage transformers’ cross-attention to correlate region and temporal features. Specifically, we use the per-segment embeddings of the proposed regions as queries in a cross-attention layer. This layer performs Multi-Head Attention over the entire sequence of spatio-temporal features from the video feature extractor, which serve as the keys and values. Consequently, this layer can fully exploit the spatio-temporal resolution to enrich the region features with relevant temporal context. Finally, we use a linear classification layer to project the enriched region features into either instrument or atomic action class probability distributions.
Our region classification head can be agnostic to the specific region proposal method and thus we experimented with multiple Mask2Former-like models ([94, 91, 12]). Nevertheless, we use our instrument segmentation baseline as the main region proposal network to obtain pixel-wise instrument instance region proposals and their per-segment embeddings. Figure 10 shows this head in the upper-right part.
@@ 4.4 Implementation Details
All versions of TAPIS involving the video feature extractor were trained on 16-frame time windows using a cosine learning rate scheduler ([57]) and an SGD optimizer, and we conducted all experiments on 4 NVIDIA Quadro RTX 800 GPUs. We adopt the following three-stage training approach, portrayed in Figure 17 of the Supplementary Material:
1. Multi-Task Long-Term Recognition: We trained the video feature extractor to perform simultaneous phase and step recognition using separate classification heads for each task and a combined cross-entropy loss. We initialized MViT with its pretrained weights from Kinetics 400 ([44]) and trained it for 30 epochs on temporal windows centered on all GraSP frames sampled at 1fps. We use a window stride of 1s, a batch size of 24, a base learning rate of $11e-$ 1, an end learning rate of $31e-$ 3, and 5 warm-up epochs.
2. Instrument Segmentation Baseline (TAPIS Frame): We independently trained Mask2Former for frame-wise instrument segmentation on the annotated keyframes using its pretrained weights on MS-COCO ([52]) with Swin Large ([55]) backbone and a batch size of 8. The training spanned 100 epochs using an ADAMW optimizer ([58]) with a base learning rate of $51e-$ 5 decayed by a factor of 0.1 at epochs 50 and 75. We selected the top 5 scoring regions with confidence scores greater than 0.1 for inference and region proposal.
3. Short-Term Task Fine-tunning: Using the pretrained weights from phase and step recognition, we fine-tuned TAPIS for instrument reclassification and atomic action detection. We froze the instrument segmentation baseline and used its precalculated instrument regions for both tasks. Both tasks were trained on all the keyframes annotated for short-term tasks as the window’s central frame, with batch size 24. For instrument classification, we trained TAPIS for 20 epochs with a window stride of 0.5 seconds, a batch size of 24, a base learning rate of $21.25e-$ 2, and an end learning rate of $41e-$ 4. For atomic action detection, training spanned 50 epochs using a window stride of 0.033s, a binary cross-entropy loss, a base learning rate of $11.25e-$ 1, and an end learning rate of $21e-$ 2.
Benchmarking: We performed all optimization and design experiments in the defined cross-validation folds. For testing, we retrained our best-performing models on the combined training set (fold one and two) and evaluated the test set to fully utilize the available training data.